OpenStudy (anonymous):

how differentiate this question (x^2-4)(x^2+3)^6

OpenStudy (diyadiya):

use product rule \[\frac{d}{dx}(uv)=v \frac{du}{dx}+u \frac{dv}{dx}\]

OpenStudy (diyadiya):

\[\frac{d}{dx}(x^2-4)(x^2+3)^6=(x^2+3)^6 \frac{d}{dx}(x^2-4)+(x^2-4) \frac{d}{dx}(x^2+3)^6\]

OpenStudy (anonymous):

= 2x (x²+3)^6 + 12x (x²-4) ( x² + 3) ^5 = 2x ( x² + 3) ^5 * [ x² + 3) + 6 (x²-4) ] = 2x ( x² + 3) ^5 * ( 7x² - 21)